2) What is the net equipment cost, given the following, when a new piece of equipment replaces an old one:

Old equipment sells for $125,000

Book value of old equipment 22,000

TaxRate 40%

New equipment cost 800,000

Site survey 18,000

Installation cost 20,000
Answer #1

Solution:-

Step 1: Calculation of proceeds from selling old equipment:

Capital gain tax on sale= (Sale consideration-book value)*40% = (125,000-22,000)*40%= $41,200

Proceeds from sale= Sale consideration- capital gain tax = $125,000-$41,200 = $83,800

Step 2: Calculation of gross cash outflow on purchase of new equipment

Gross cash outflow= New equipment cost + Site survey + Installation cost

Gross cash outflow= 800,000 + 18,000 + 20,000 = $838,000

Step 3: Calculation of net equipment cost

Net equipment cost= Gross cash outflow for new equipment - proceeds from sale of old equipment

Net equipment cost= $838,000 - $83,800 = $754,200

Thus, the net equipment cost is $754,200 for the buyer.

Please feel free to ask any doubts in the comments section.